OpenStudy (anj123):

Find a formula in terms of k for the entries of A^k, where A is the diagonalizable matrix below, and the eigenvalues of A are −1 and 2. A = -10 18 -6 11

OpenStudy (anj123):

P = 3 2 2 1 P^-1 = -1 2 2 -3 Then what?

OpenStudy (empty):

Now you have the eigenvalues, so depending on how you set up P and P inverse you'll have this: \[A=P^{-1}DP\] So any power of A, such as \(A^2\) will be \[A^2=P^{-1}DPP^{-1}DP=P^{-1}D^2P\] Notice on the inside there we have P and its inverse, so we just have the identity! \(PP^{-1}=I\) This will work for larger powers just as well, \[A^k=P^{-1}D^k P\] What are the entries of D? They're the eigenvalues of A, check this to make sure, and then you can see what happens when you raise a diagonal matrix to a power.

OpenStudy (anj123):

so D = -1 0 0 -1 or is it D = 2 0 0 2

OpenStudy (empty):

Neither, it will be either diag(-1, 2) or diag(2, -1) depending on which eigenvalue is which, really you're writing the eigenvalue equation in this new way: \[A v_1 = \lambda_1 v_1\]\[A v_2 = \lambda_2v_2\] Combine them so that these column vectors fill a 2x2 matrix: (work out the matrix multiplication and try to understand as best you can, tell me where/if you get stuck:) \[A\begin{bmatrix} v_1 & v_2 \end{bmatrix} = \begin{bmatrix} v_1 & v_2 \end{bmatrix}\begin{pmatrix} \lambda_1 &0 \\0 &\lambda_2 \end{pmatrix}\] This matrix of eigenvectors are our P or \(P^{-1}\) matrix (depending on what you call P) and can be inverted: \[A\begin{bmatrix} v_1 & v_2 \end{bmatrix} = \begin{bmatrix} v_1 & v_2 \end{bmatrix}\begin{pmatrix} \lambda_1 &0 \\0 &\lambda_2 \end{pmatrix}\] Multiply on the right by the inverse, \[A\begin{bmatrix} v_1 & v_2 \end{bmatrix}\begin{bmatrix} v_1 & v_2 \end{bmatrix}^{-1} = \begin{bmatrix} v_1 & v_2 \end{bmatrix}\begin{pmatrix} \lambda_1 &0 \\0 &\lambda_2 \end{pmatrix}\begin{bmatrix} v_1 & v_2 \end{bmatrix}^{-1}\] \[A= \begin{bmatrix} v_1 & v_2 \end{bmatrix}\begin{pmatrix} \lambda_1 &0 \\0 &\lambda_2 \end{pmatrix}\begin{bmatrix} v_1 & v_2 \end{bmatrix}^{-1}\]

OpenStudy (anj123):

So, basically in order to find D, you have to follow this formula each time D =P^-1AP

OpenStudy (irishboy123):

.

OpenStudy (empty):

Yeah, basically, but it's really just lumping all the eigenvalue/eigenvector formulas of the form: \[Ax_1 = \lambda_1x_1\]\[Ax_2 = \lambda_2x_2\]\[\cdots\] into just one matrix equation where the matrix P has all the eigenvectors and the diagonal has all the eigenvalues \[AP=PD\]